Project

General

Profile

20170131 » History » Version 20

Nebrensky, Henry, 01 February 2017 08:49
DS

1 1 Boyd, Steven
h1. Ops/Comm Scheduling meeting
2
3 2 Boyd, Steven
Location : G06 in R66 at 2pm
4 1 Boyd, Steven
5
Phone access :  ReadyTalk number 
6
US 866-740-1260
7
UK 0800-496-0576
8
Swiss 0800-70-53-51
9
Italy 800182592
10
Access Code: 2155402
11
12
* SS Magnet status
13
* Hydrogen status
14
* Issues arising from previous running
15 4 Boyd, Steven
** [[Hall_Task_List_Dec16Feb17|Shutdown work in Jan and Feb]]: "present version":https://micewww.pp.rl.ac.uk/attachments/download/8158
16 1 Boyd, Steven
* Operations :
17 6 Boyd, Steven
** C&M System Check report (Ajit/Ed)
18 1 Boyd, Steven
** Establishing a timetable for the runup : when should systems (including magnets) be ready for running?
19
*** Beam schedule
20
**** Run up on Jan 30 - Feb 12.
21
**** Start of user cycle : Feb 14
22
**** Start of data-taking : Feb 15
23 14 Boyd, Steven
*** Reserve shifter support from Feb 10 - Feb 14 
24 1 Boyd, Steven
*** MOM formally starts on Feb 9 (but should be present for the runup activities but will not be on 24-hour call)
25
*** Possible runup plan (framework for discussion)
26 7 Boyd, Steven
**** Monday 6th : Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running
27 15 Boyd, Steven
**** Tuesday 7th (subject to meeting on Monday) : Bring up magnets to running configuration, take some pulser data and take magnets down
28 5 Boyd, Steven
**** Saturday 11th : Expert led run without solenoids + Beam tune
29
**** Monday 13th : Triage actions arising from expert led run
30 8 Boyd, Steven
**** Tuesday 14th : Bring up magnets in the evening and monitor overnight by reserve shift.
31 16 Boyd, Steven
**** Wednesday 15th : Data-taking start at 08:00 
32 1 Boyd, Steven
33
** Shift signup for first half of Feb is viable now (but not full)
34
35 10 Boyd, Steven
| *Shift Block ID*   | # signed up  |  # shifters still needed  | 
36 1 Boyd, Steven
|         1          |     6        |            0              |
37 11 Boyd, Steven
|         2          |     4        |            2              | *Need one overnight shift per night (5 nights) for monitoring* |
38
|         3          |     4        |            2              | *Need one overnight shift per night (5 nights) for monitoring* |
39 1 Boyd, Steven
|         4          |     4        |            1              |
40
41 12 Boyd, Steven
*Magnet monitors do not need to be trained shifters*
42
43 13 Boyd, Steven
** BLOCs are needed : Feb 20 - 24 and Feb 27 - Mar 1
44 1 Boyd, Steven
45
* AOB:
46
47
* Last meeting : [[20170124|Tuesday Jan 24 2017]]
48
* Next meeting : [[20170206|Tuesday Feb 6 2017]]
49 17 Nebrensky, Henry
50
h2. Minutes
51
52
* CC Magnet status (J Boehm)
53
  All 3 CC magnets are now in flip mode; cold and full of helium. Everything is presently being run into shorts for the CC IOC testing.
54
  Handover from Pierrick is in progress. A few GUIs have been changed.
55
  Are seeing instability in SSU power supply (into shorts): last week suddenly ramped from 150 to 215 A; lost control of GUI; ramped down manually.
56
  Similar last night, with control cables disconnected from the AMI controller - again unstable and major changes in current on its own, implies a hardware issue.
57
  Now running with a different controller - seeing issues still. May possibly be associated with a purely resistive load.  
58
59
* Hydrogen status (M Tucker)
60
  Cooled down in R9 with Neon, made 2 to 2.5 litres of liquid, less than 1 K difference between condenser and absorber.
61
  Expect about 1 l every 6 hours, so over 5 days to fill with LH2 needing a 24hour rota. 
62
63
* Issues arising from previous running (DC)
64
  Need a comprehensive list of encountered issues, else can't assume they will be addressed before running resumes. We are now well past halfway through the shutdown.
65
* Shutdown work in Jan and Feb ( "official version":https://micewww.pp.rl.ac.uk/attachments/download/8158 )
66
  Issues/Progress on workplan items (using a stale list):
67
68
  ** PT100s - done
69
  ** ODH sensor service - done
70
  ** POMPOMs sensor test - done
71
  ** Neutron Monitor: ISIS can't do it; will return to MICE for us to make arrangements. Continue with temporary monitor this cycle.
72
  ** Proton Absorbers - remedial work done
73
  ** "New" chiller plumbing - done
74
  ** #1878 Trench heat exchanger to counter-flow - done, possibly some small leaks
75
76
  ** Diffuser iris 4: needs next SSU run: Tuesday 7th
77
     Compressed air: Activair coming to check the outside compressor to see why it didn't come on.
78
79
  ** #1399 MICEnet lockout - tomorrow morning
80
81
  ** #1888 Diffuser air - reported done
82
  ** Airflow monitoring changes - done
83
  ** PPS Functional Test - done
84
  ** Earth leakage tests - done
85
  ** New @target1ctl@ - issues with O/S and libraries...
86
  ** R78 - replacement flanges in progress
87
88 20 Nebrensky, Henry
* The Decay Solenoid is cold. Only known issue is the episodes of fridge instability seen last run.
89
  Presently the cooling lead flow is on, to check the stability of the fridge; otherwise ready to power.
90
91 17 Nebrensky, Henry
* C&M System Check report (Ajit/Ed)
92
  CC Magnet UIs/IOCs tested last week - changes have been made on various machines, need to be aggregated.
93
  Target SM tested today.
94
  Need to test Run Control: Thursday?
95
  Need to review the tags listed in the CDB
96
  The NIKHEF Hall probes have been separated out into a different IOC.
97
  Most of the hardware stuff is done. 
98
99
* Run-up plan 
100
** Roof water/chiller needs a load test with the air-blast active - tomorrow or Thursday subject to Hall work and state of pipes 
101
** Monday 6th: Preparation meeting with Ed/Ajit/Colin/Josef/Others to go over readiness for running; remove shorts from link boxes, ATW and other prep.
102
** Tuesday 7th: Bring up CC magnets to running configuration, take some pulser data and take magnets down; _check Diffuser operation_
103
** Saturday 11th - Half-day of ISIS beamtime: expert led run without CC Magnets + Beambump tune (Hayley, NB Ed has to leave mid-afternoon)
104
** Monday 13th: Triage actions arising from expert led run; BLOC training on everything that isn't target
105
** Tuesday 14th: Bring up magnets in the evening and soak-test overnight (reserve shifters).
106
** Wednesday 15th: Data-taking start at 08:00 
107
108
* Shift signup:
109 18 Nebrensky, Henry
  With the LH2 install going ahead there will be no running during second half of 2016/05 and thus no available shift allocation - MICE that have not yet made shift quota take note. 
110 17 Nebrensky, Henry
111
* AOB:
112
  Linde Gas Analyser webcam - where has it gone? It is being PAT tested.
113
114
h3. Actions
115
116
* JBoehm to get SS operating info from PHanlet (ongoing)
117
* JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
118
* DC to collate list of issues arising from previous running 
119
* Status of @miceraid5@ disk cleanup (stands)
120
* -AKurup to make and circulate plan for NIKHEF probe IOC-
121
* -SBoyd - ISIS machine-physics beam time-
122
* DC to ensure Health Physics have checked temporary neutron monitor (stands)
123
* SBoyd - ensure MICE-wide discussion of MOM C&M panel
124 19 Nebrensky, Henry
* Henry to get gaps in BLOC rota plugged